OpenStudy (anonymous):

Simplify and give your answer in the form f(x)/g(x): [(6/(x-1))-(1/(x+1))]/[(x/(x-1))+(1/(x+1))]

OpenStudy (mathteacher1729):

How far have you gotten with this problem?

OpenStudy (anonymous):

Not very.

OpenStudy (anonymous):

multiply top and bottom (carefully) by \[(x-1)(x+1)\] to clear the fraction. that is probably easiest

OpenStudy (mathteacher1729):

It's best to break this into pieces. Start with just the stuff in the numerator. Simplify that. Then do the same with the denom.

OpenStudy (anonymous):

you will get \[6(x+1)-(x-1)\] in the numerator and \[x(x+1)+(x-1)\] in the denominator. then so some algebra to multiply out and combine like terms. should be good from there

OpenStudy (anonymous):

careful with the distributive law and so one, but now you have \[\frac{6(x+1)-(x-1)}{x(x+1)+x-1}\] so that is most of the work

OpenStudy (anonymous):

why do you multiply x(x+1) in the denom?

OpenStudy (anonymous):

ok you have the denominator of \[\frac{x}{x-1}+\frac{1}{x+1}\] yes?

OpenStudy (anonymous):

unless i read it incorrectly. i mean this is your entire denominator.

OpenStudy (anonymous):

so now when you multiply by \[(x+1)(x-1)\] you get \[(\frac{x}{x-1}+\frac{1}{x+1})(x+1)(x-1)\]

OpenStudy (anonymous):

in the first term the x - 1 cancel giving \[x(x+1)\] and in the second term the x +1 cancel giving \[1(x-1)\] so the whole denominator becomes \[x(x+1)+x-1\]

OpenStudy (anonymous):

of course you should then write this as \[x^2+2x-1\]

OpenStudy (anonymous):

ok. I get it now. You've been a great help.